A 0.300-kg piece of iron initially at 315 ºC is dropped into 0.400 L of water initially at 530ºR. Assuming that all heat transfer occurs between the iron and the water, calculate the final temperature.


1
Expert's answer
2021-01-18T01:35:52-0500

Heat change of iron = Heat change of Water .


0.3 × 1000 × 0.450 × ( 273 +315 - T )


= 400 × 4.18 × ( T - ( 530 + 273 ) )


300 × 0.450 × ( 588 -T )

= 400 × 0.450 × ( T - 803 ) .


135 ( 588 -T ) = 180 ( T - 803 ) .


135 × 588 - 135 T = 180 T - 180 × 803


180 × 803 - 135 × 588 = 180 T - 135 T


144540 - 79380 = 45 T .


65160 = 45 T .


T = 65160 / 45 = 1448 K .


Temperature in degree celcius = ( 1448 - 273) = 1175 degree .